### v2

# Maximum Inner Product is Query-Scaled Nearest Neighbor

Abstract.
Maximum Inner Product Search ($\mathsf{MIPS}$) for high-dimensional vectors is pivotal across databases, information retrieval, and artificial intelligence. Existing methods either reduce $\mathsf{MIPS}$ to Nearest Neighbor Search ($\mathsf{NNS}$) while suffering from harmful vector space transformations, or attempt to tackle $\mathsf{MIPS}$ directly but struggle to mitigate redundant computations due to the absence of the triangle inequality. This paper presents a novel theoretical framework that equates $\mathsf{MIPS}$ with $\mathsf{NNS}$ without requiring space transformation, thereby allowing us to leverage advanced graph-based indices for $\mathsf{NNS}$ and efficient edge pruning strategies, significantly reducing unnecessary computations. Despite a strong baseline set by our theoretical analysis, we identify and address two persistent challenges to further refine our method: the introduction of the Proximity Graph with Spherical Pathway ($\mathsf{PSP}$), designed to mitigate the issue of $\mathsf{MIPS}$ solutions clustering around large-norm vectors, and the implementation of Adaptive Early Termination (AET), which efficiently curtails the excessive exploration once an accuracy bottleneck is reached. Extensive experiments reveal that our method is superior to existing state-of-the-art techniques in search efficiency, scalability, and practical applicability. Compared with state-of-the-art graph-based methods, it achieves an average 35% speed-up in query processing and a 3$\times$ reduction in index size. Notably, our approach has been validated and deployed in the search engines of Shopee, a well-known online shopping platform. Our code and an industrial-scale dataset for offline evaluation will also be released to address the absence of e-commerce data in public benchmarks.
PVLDB Reference Format:
PVLDB, 18(6): 1770 - 1783, 2025.
doi:10.14778/3725688.3725705 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 6 ISSN 2150-8097.
doi:10.14778/3725688.3725705
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/ZJU-DAILY/PSP.
1. Introduction
Maximum Inner Product Search ($\mathsf{MIPS}$) is essential across various artificial intelligence and information retrieval applications (Asai et al., 2023; Zhou et al., 2022; Seo et al., 2019; Lewis et al., 2020). The demand for managing large-scale, high-dimensional data — fueled by advancements of large language models (Zhao et al., 2023b) and retrieval-augmented generation (Asai et al., 2023) — has garnered significant attention within the database community (Guo et al., 2022; Pan et al., 2023; Yang et al., 2020). However, efficient accurate $\mathsf{MIPS}$ remains a formidable challenge (Teflioudi et al., 2015; Indyk and Motwani, 1998; Weber et al., 1998; Chen, 2020). In response, there has been a shift towards approximate $\mathsf{MIPS}$, which trades minimum accuracy for substantial gain in speed.
For the approximate $\mathsf{MIPS}$ problem, two primary paradigms have emerged. The first focuses on the Inner Product (IP) metric, establishing specialized theoretical frameworks to tackle $\mathsf{MIPS}$ challenges (Morozov and Babenko, 2018; Liu et al., 2020; Tan et al., 2021; Guo et al., 2016; Zhang et al., 2023a; Guo et al., 2020; Bruch et al., 2023). However, the absence of triangle inequality in the IP space hinders them from efficiently reducing redundant computations (Tan et al., 2021), particularly in recent promising graph-based methods (Morozov and Babenko, 2018; Liu et al., 2020). This deficiency impacts query performance and increases memory usage, as these methods lack theory foundations to prune edges as effectively as advanced $\mathsf{NNS}$ graphs (Malkov and Yashunin, 2018; Fu et al., 2021, 2019). Specifically, the widely-used greedy graph search algorithm (Algorithm 1) necessitates checking all neighbors at each step to approach the query more closely during traversal, often leading to unnecessary checks that degrade efficiency (Wang et al., 2021b; Liu et al., 2020; Tan et al., 2021).
The second paradigm addresses the $\mathsf{MIPS}$ problem by reformulating it as an $\mathsf{NNS}$ problem, enabling the use of established $\mathsf{NNS}$ methods (Shrivastava and Li, 2014, 2015; Neyshabur and Srebro, 2015; Li et al., 2018; Yan et al., 2018; Zhao et al., 2023a; Zhou et al., 2019) that exploit the triangle inequality for efficiency. These methods involve various space transformations, which, while effective to some extent, often rely on strong theoretical assumptions (Zhou et al., 2019), can introduce potential structural distortions (Zhao et al., 2023a; Morozov and Babenko, 2018), and exhibit inefficiencies in data updating (elaborated in §2). These significantly restrict their practical applicability.
Our research explores a fundamental question: Can we leverage the efficient computation reduction properties without altering the vector space? We begin with an intriguing observation derived from comparing the search routing behaviors of $\mathsf{MIPS}$ and $\mathsf{NNS}$ on the same Euclidean proximity graph, such as an MSNET (Fu et al., 2019). Specifically, using Algorithm 1, the search iteratively moves from a node to its neighbor, minimizing the distance to the query under given metrics. As shown in Figure 1, the search paths under $\mathsf{MIPS}$ and NNS initially overlap. The main discrepancy arises when $\mathsf{MIPS}$ moves beyond the query into outer spherical areas of larger norms. According to the homogeneity of the IP metric, i.e., $\langle\mu q,p\rangle=\mu\langle q,p\rangle$, their paths and solutions coincide when a scalar scales up $q$ until its norm exceeds the $\mathsf{MIPS}$ solution.
Figure 1. An illustration of the overlaps and discrepancies between the paths of NNS and MIPS on a Euclidean proximity graph. The overlaps hold mainly before the norms of explored nodes exceed the norm of the query.
We theoretically verify the above observations by showing that $\mathsf{MIPS}$ for query $q$ can be equated to $\mathsf{NNS}$ for a scaled query $q^{\prime}=\mu q$ on an MSNET, via introducing a proper scalar $\mu$. This equality enables using advanced $\mathsf{NNS}$ graph indices for $\mathsf{MIPS}$ without altering the vector space, thus avoiding the associated drawbacks. Additionally, this breakthrough lays the groundwork for the first theoretical analysis of search complexity in the graph-based $\mathsf{MIPS}$ context. Our experiments on real-world datasets further validate this theory (§5). Despite these advancements, two practical challenges persist: the tendency of $\mathsf{MIPS}$ solutions to congregate in large norm regions and the excessive exploration that occurs once a precision bottleneck is reached. We propose an innovative index, Proximity Graph with Spherical Pathway ($\mathsf{PSP}$), to handle the biased distribution of $\mathsf{MIPS}$ solutions. Additionally, we introduce a novel Adaptive Early Termination (AET) mechanism to mitigate excessive explorations.
Contributions. Our contributions are highlighted as follows:
(1) Theoretical Foundation: This work pioneers in establishing comprehensive theoretical foundations for graph-based $\mathsf{MIPS}$.
(2) Scalable Solution: We introduce $\mathsf{PSP}$, a scalable graph index with spherical pathways and adaptive search initialization for the biased distribution of $\mathsf{MIPS}$ solutions. We propose an adaptive early termination mechanism to mitigate the excessive exploration.
(3) Extensive Experiments on 8 real-world datasets validate PSP’s theoretical soundness, efficiency, and scalability, highlighting an average 35% speed-up in query process and a 3$\times$ reduction in index size over leading graph methods. PSP is the only one reaching 95% recall under 5ms and thus has been deployed in the large-scale search engines of Shopee, which showcases the application viability.
(4) New Industry-scale Dataset: We release a new dataset derived from industry-scale traffic data collected from the Shopee search scenario. Our dataset fills the absence of e-commercial modality in public benchmarks, facilitating research in $\mathsf{MIPS}$ and $\mathsf{NNS}$.
Roadmap. §2 introduces the background of the $\mathsf{MIPS}$ problem and two main paradigms; §3 presents our theoretical foundations; §4 introduces practical solutions ($\mathsf{PSP}$ and AET); §5 outlines the research questions and provides a comprehensive analysis based on the experimental outcomes; §6 discusses the new dataset and limitations; §7 reviews previous related works for broader interests and §8 concludes this paper.
2. Preliminaries and Background
[…]
The Inner Product (IP) serves as a fundamental and effective similarity metric, widely used in artificial intelligence and machine learning (Asai et al., 2023; Radford et al., 2021a). With the proliferation of data represented and stored in high-dimensional vectors (Guo et al., 2022; Pan et al., 2023; Zhao et al., 2023b), MIPS has become increasingly important (Zhao et al., 2023a; Guo et al., 2020; Morozov and Babenko, 2018). MIPS is formally defined as:
Definition 1 (Maximum Inner Product Search (MIPS)).
Given a query $q\in\mathbb{R}^{d}$, and a dataset $\mathcal{D}\subset\mathbb{R}^{d}$, the $\mathsf{MIPS}$ problem aims to find a vector $p\in\mathcal{D}$ that maximizes the inner product with $q$:
(1) $p=\operatorname*{arg\,max}_{p\in\mathcal{D}}\langle p,q\rangle$
Several approaches have been proposed to solve the MIPS problem in sublinear time (Abuzaid et al., 2019; Teflioudi and Gemulla, 2016; Li et al., 2017). However, they suffer from inefficiencies in query process. Consequently, researchers have explored approximate $\mathsf{MIPS}$ methods, which trade a minimum accuracy loss for significantly improved retrieval efficiency. Formally, we have:
Definition 2 ($\epsilon$-Maximum Inner Product Search ($\epsilon$MIPS)).
Given a query $q\in\mathbb{R}^{d}$, a dataset $\mathcal{D}\subset\mathbb{R}^{d}$, and an approximation ratio $\epsilon\in(0,1)$, let $p^{*}\in\mathcal{D}$ be the $\mathsf{MIPS}$ solution of q, the $\epsilon$$\mathsf{MIPS}$ problem aims to find a vector $p\in\mathcal{D}$ satisfying $\langle p,q\rangle\geq\epsilon\cdot\langle p^{*},q\rangle$.
While defining $k$-$\mathsf{MIPS}$ and $k$-$\epsilon$$\mathsf{MIPS}$ to find the top $k$ solutions is straightforward, we omit these details here for brevity. Note that the main difference between MIPS and a similar field, Nearest Neighbor Search (NNS), lies in the metric used (Pan et al., 2023) (inner product and Euclidean distance respectively) Recent efforts aim to address both $\epsilon$NNS and $\epsilon$$\mathsf{MIPS}$ problems with similar strategies: reducing the cost of distance calculations and pruning the search space (Zhao et al., 2023a; Ghojogh et al., 2018; Guo et al., 2020; Zheng et al., 2016). To better motivate our study, we revisit the $\epsilon$$\mathsf{MIPS}$ techniques from a novel view: whether transforming the vector space.
Non-transformation-based methods use the Inner Product (IP) for both index structures and search algorithms (Morozov and Babenko, 2018; Liu et al., 2020; Tan et al., 2021; Guo et al., 2016; Zhang et al., 2023a; Guo et al., 2020). While these methods preserve the integrity of information in the original space, they often suffer from large indices and inferior search performance due to their inability to reduce redundant computations effectively. This inefficiency primarily stems from the absence of properties analogous to the triangle inequality in the Euclidean metric. In contrast, the triangle inequality significantly reduces the search space in $\epsilon$NNS under the Euclidean metric (Wang, 2011).
Transformation-based methods, inspired by the successes in NNS, attempt to reframe the MIPS problem into an NNS problem through two typical space transformation techniques:
Möbius transformation (Zhou et al., 2019) aims to establish an isomorphism between the Delaunay graph for the IP metric (Morozov and Babenko, 2018) and that for the Euclidean metric (Malkov et al., 2014) by scaling the vectors:
Definition 3 (Möbius Transformation).
Given a database $\mathcal{D}\subset\mathbb{R}^{d}\backslash\{0\}$, the möbius transformation modifies $p\in\mathcal{D}$ by $p/\lVert p\rVert^{2}$.
The isomorphism established by Möbius transformation comes with strong assumptions, such as requiring the origin point to be within $\mathcal{D}$’s convex hull and the data to be independently and identically distributed (IID), which are often violated in practice. The origin-contained assumption could easily be violated with one-hot or multi-hot representations. Verifying that the origin is not contained in such convex hull is straightforward. Additionally, the IID assumption is often violated because the representations generated by ML models are usually correlated in different dimensions (Radford et al., 2021a; Schuhmann et al., 2022).
XBOX transformation (Zhao et al., 2023a) adopts an asymmetric approach by elevating both the dataset and query vectors to higher dimensions:
Definition 4 (XBOX Transformation).
Given a database $\mathcal{D}\subset\mathbb{R}^{d}$, and a query $q$. The XBOX transformation maps $\forall p\in\mathcal{D}$ to $p^{\prime}=\left[p;\sqrt{M^{2}-\lVert p\rVert^{2}}\right]$, and maps $\forall q\in\mathcal{R}^{d}$ to $q^{\prime}=\left[q;0\right]$: where [;][;][ ; ] denotes vector concatenation and $M\geq\max_{p\in\mathcal{D}}||p||$.
Thus, we have $\langle p,q\rangle=\frac{1}{2}\left(\lVert q\rVert^{2}+M^{2}-\delta(p^{\prime},q^{\prime})\right)$, a conversion from IP to Euclidean with constant offset. Although this transformation is widely used (Neyshabur and Srebro, 2015; Keivani et al., 2018; Pham, 2021; Yan et al., 2018; Huang et al., 2018), it has significant drawbacks like potential structural distortions (Morozov and Babenko, 2018) and difficulties in determining an appropriate $M$: A large $M$ may reduce the distinguishability of different points because it contributes significantly to $\delta(P(p),Q(q))$. Conversely, a small $M$ may cause inflexibility in data updates when the new points’ norms exceed $M$.
Graph-based methods for $\mathsf{MIPS}$ is the focus of this paper. While previous approaches vary in their graph index structures, they predominantly employ a similar search algorithm, often referred to as a greedy walk (Algorithm 1). This algorithm iteratively moves to a neighbor of the current node that is closer to the query, with the search complexity dominated by the walking steps and the graph’s average out-degree (Fu et al., 2019). Notably, non-transformation graph-based methods have struggled with implementing efficient edge pruning strategies due to the absence of the triangle inequality, leading to graphs with high average degrees and subsequently increased search complexity. In contrast, advanced graph-based $\mathsf{NNS}$ methods (Malkov and Yashunin, 2018; Fu et al., 2021, 2019) can dramatically reduce edge quantity while maintaining graph connectivity and short search paths, significantly enhancing efficiency. In the following section, this paper will explore how to leverage these properties for the $\mathsf{MIPS}$ problem by providing theoretical foundations to align graph-based $\mathsf{MIPS}$ with $\mathsf{NNS}$ without vector space transformations.
Input: Dataset $\mathcal{D}$, Graph $G$, query $q$, candidate set size $l_{s}$, result set size $k$.
Output: Top $k$ result set $R$.
1 $R\leftarrow\emptyset$; $q^{\prime}\leftarrow\mu q$; $Q\leftarrow\emptyset$;
2 $P\leftarrow$ random sample $l_{s}$ nodes from $G$;
3 for each node $p$ in $P$ do
4       $Q$.add($p$,$\delta(p,q^{\prime})$);
[…]
6 while $Q$.size() do
7       $p\leftarrow Q.pop()[0]$; $R.insert((p,\delta(p,q^{\prime})))$
8       if visited($p$) then
9            continue;
10      $N_{p}\leftarrow$ neighbors of $p$ in $G$
11       $Q\leftarrow$ batch_insert($Q,N_{p}$)
12       $Q.resize(l_{s})$;$R.resize(k)$triangleright delete larger-distance nodes
13
14
return $R$
Algorithm 1 Greedy Search For Graphs(Prokhorenkova and Shekhovtsov, 2020)
3. Theoretical Foundations
This section aims to equate $\mathsf{MIPS}$ and $\mathsf{NNS}$ theoretically on Euclidean proximity graphs. This alignment allows us to capitalize on the beneficial features of established graph-based $\mathsf{NNS}$ methods, such as efficient edge pruning, strong connectivity, and robust theoretical guarantees, to enhance graph-based $\mathsf{MIPS}$ and analyze its search behavior. Formally, a proximity graph is defined as:
Definition 5 (Proximity Graph).
Given a dataset $\mathcal{D}$, a proximity graph $G$ on $\mathcal{D}$ can be denoted by $G=(V,E)$, where $V$ is the node set, and $E$ is the edge set. Each node $v_{i}\in V$ represents a vector $p_{i}\in\mathcal{D}$, while $(v_{i},v_{j})\in E$ if and only if $(v_{i},v_{j})$ satisfies a given criteria.
Within our theoretical alignment, the specific type of proximity graph—whether a Delaunay graph (Aurenhammer, 1991), a Navigating Small World (NSW) graph (Malkov and Yashunin, 2018), or a Monotonic Relative Neighborhood Graph (MRNG) (Fu et al., 2021)—is not fixed. The choice of graph does not affect the validity of our theory. Notably, part of our theory also is not limited to graph-based context (Theorem 1) and can be applied to any type of $\mathsf{MIPS}$ methods, such as hashing and quantization.
3.1. Equivalence under Scaled Query
First, we establish a mapping $q^{\prime}=f(q)$, such that (1) the $\mathsf{MIPS}$ solutions for $q^{\prime}$ align with those for $q$, and (2) the $\mathsf{NNS}$ solutions for $q^{\prime}$ align with the $\mathsf{MIPS}$ solutions for $q$. This mapping enables the use of efficient NNS algorithms to solve the $\mathsf{MIPS}$ problem without changing the origin vector space. We leverage the homogeneity of IP metric, as formalized in the following lemma:
Lemma 1.
Given a vector database $\mathcal{D}\subset\mathbb{R}^{d}$, and a scalar $\mu>0$, the $\mathsf{MIPS}$ solution for $q$ is identical to those for $q^{\prime}=\mu q$ within $\mathcal{D}$.
Proof.
Let $p^{*}\in\mathcal{D}$ be a MIPS solutions for $q$, satisfying $\forall p\in\mathcal{D}\backslash\{p^{*}\},\langle p^{*},q\rangle\geq\langle p,q\rangle$. Scaling $q$ by $\mu>0$, we have:
(2) $\langle p^{*},\mu q\rangle=\mu\langle p^{*},q\rangle\geq\mu\langle p,q\rangle=\langle p,\mu q\rangle,\forall p\in\mathcal{D}\backslash\{p^{*}\}$
Therefore, $q^{\prime}=\mu q$ and $q$ share the same $\mathsf{MIPS}$ solutions. ∎
Lemma 1 reveals that non-negative constant scaling of the query vector preserves $\mathsf{MIPS}$ solutions. Next, we will align $\mathsf{MIPS}$ and $\mathsf{NNS}$ with this scaling mapping by identifying solution space for $\mu$.
Theorem 1.
Given a vector database $\mathcal{D}\subset\mathbb{R}^{d}$ and $\forall q\in\mathbb{R}^{d}$, there exists a scalar $\bar{\mu}$ such that for $\forall\mu>\max(\bar{\mu},0)$, the nearest neighbor of $q^{\prime}=\mu q$ in $\mathcal{D}$ aligns with the $\mathsf{MIPS}$ solution for $q$.
Proof.
We exclude cases where the zero-vector or multiple identical vectors in $\mathcal{D}$ complicate the analysis. Then, we break down the proof step by step.
Case 1: Assume there is a unique $\mathsf{MIPS}$ solution $p^{*}\in\mathcal{D}$ for $q$. By Lemma 1, $p^{*}$ is also a $\mathsf{MIPS}$ solution for $q^{\prime},\forall\mu>0$.
To ensure $p^{*}$ is a nearest neighbor of $q^{\prime}=\mu q$, we need to solve:
$\lVert p^{*}-q^{\prime}\rVert<\lVert p-q^{\prime}\rVert,\quad\forall p\in\mathcal{D},p\neq p^{*}\$
By simplifying and rearranging, we find the condition for $\mu$:
(3) $\mu>\frac{\lVert p^{*}\rVert{{}^{2}}-\lVert p\rVert^{2}}{2(\langle p^{*},q\rangle-\langle p,q\rangle)},\forall p\in\mathcal{D},p\neq p^{*}$
Let $\bar{\mu}=\sup\left(\left\{\frac{\lVert p^{*}\rVert{{}^{2}}-\lVert p\rVert^{2}}{2(\langle p^{*},q\rangle-\langle p,q\rangle)}\Big{|}\forall p\in\mathcal{D},p\neq p^{*}\right\}\right)$. This space is bounded because $\mathcal{D}$ is a finite set; thus, $\bar{\mu}$ exists. Unifying conditions for $\mu$, $p^{*}$ is a nearest neighbor of $q^{\prime}$ when $\mu>\max(\bar{\mu},0)$.
Case 2: Considering there may exist multiple $\mathsf{MIPS}$ solutions for $q\in\mathcal{D}$, let $\mathcal{P}^{*}=\left\{p^{*}|\langle p^{*},q\rangle\geq\langle p,q\rangle,\forall p\in\mathcal{D},p\neq p^{*}\right\}$ denote this solution set. We can find a $p_{n}^{*}\in\mathcal{P}^{*}$ such that $p_{n}^{*}$ is the nearest neighbor of $q^{\prime}=\mu q$ with an appropriate $\mu$. By following a similar procedure in Case 1, we can get the solution space for $\mu$ as:
$\mu>\max(\bar{\mu},0)=\max(\sup\left(\left\{\frac{\lVert p^{*}\rVert{{}^{2}}-\lVert p\rVert^{2}}{2(\langle p^{*},q\rangle-2\langle p,q\rangle)}|p\in\mathcal{D}\setminus\mathcal{P}^{*}\right\}\right),0)$
Summarizing Case 1 and 2, we can identify a scalar boundary $\bar{\mu}$ such that $\forall\mu>\max(\bar{\mu},0)$, there exists a point $p^{*}\in\mathcal{D}$ which is the nearest neighbor of $q^{\prime}=\mu q$ and also a $\mathsf{MIPS}$ solution of $q$. ∎
By Theorem 1, we establish the existence of a scalar $\mu$ and a corresponding node $p^{*}$ that unifies the solutions for the $\mathsf{NNS}$ and $\mathsf{MIPS}$ problems for a scaled query $q^{\prime}=\mu q$. This identification alone does not ensure we can retrieve $p^{*}$ under the IP metric as efficiently as $\mathsf{NNS}$ algorithms on a Euclidean proximity graph. To address this, we aim to demonstrate the equivalence of search behavior under IP and Euclidean metrics within the same graph structure:
Theorem 2.
Given a proximity graph $G=(V,E)$ and a query $q\in\mathbb{R}^{d}$, when using the standard Graph Nearest Neighbor Search ($\mathsf{GNNS}$(Prokhorenkova and Shekhovtsov, 2020) algorithm to decide the $\mathsf{MIPS}$ solution for $q$, there exists a scalar $\bar{\mu}$ such that for $\forall\mu>\max(\bar{\mu},0)$, we can identify a node $p^{*}\in\mathcal{N}_{o}$ that satisfies: $p^{*}\in\left\{\operatorname*{arg\,max}_{p\in\mathcal{N}_{o}}\langle p,q\rangle\right\}\cap\left\{\operatorname*{arg\,min}_{p\in\mathcal{N}_{o}}\delta(p,q^{\prime})\right\}$. Here, $o$ can be any node on the search path of $\mathsf{GNNS}$ regarding the query $q^{\prime}=\mu q$, and $\mathcal{N}_{o}=\left\{p|(p,o)\in E\right\}$ denotes $o$’s neighbor nodes.
Proof Sketch.
To align the search paths under both metrics as per Algorithm 1, we can unify the node-selection behavior by localizing the problem: solving for the optimal $p^{*}$ among the neighbors at each greedy step, akin to the conditions established in Theorem 1. Let $l$ be the number of steps in the path. Define $U=\left\{\bar{\mu}_{i}|1\leq i\leq l\right\}$. it’s not hard to verify that there exists a scalar boundary $\sup\left(U\cup\{0\}\right)$ forces the Algorithm 1 under both IP and Euclidean metrics to generate the same search path targeting query $\mu q$. Please refer to (Chen et al., 2025) for the complete proof. ∎
Remarks. From Theorem 1 and Theorem 2, we discern a compelling duality between the $\mathsf{MIPS}$ and $\mathsf{NNS}$ paradigms. These findings suggest that by appropriately scaling the query $q$, we can modulate the overlap in their search paths on a Euclidean proximity graph. By choosing an appropriate $\mu$, the search behavior becomes invariant to the metric used. Our experimental validations further corroborate this theoretical insight on real-world datasets (§3).
Notably, in practice, it is unnecessary to find a specific $\mu$ to tackle the $\mathsf{MIPS}$ problem (Lemma 1). Theorems 1 and 2 demonstrate that the IP metric can be directly applied on a Euclidean proximity graph to perform IP-greedy search with Algorithm 1.
3.2. Efficiency Analyses for Graph-Based MIPS
While our theoretical framework permits the use of any Euclidean Proximity Graph, only a few, such as MRNG (Fu et al., 2019) and SSG (Fu et al., 2021), come with robust theoretical guarantees. To overcome the high-degree issue prevalent in non-transformation graph-based methods (Morozov and Babenko, 2018; Liu et al., 2020), we aim to identify a sparse Euclidean graph that ensures a low amortized search complexity. In this paper, we focus on SSG for its sparsity and additional theoretical guarantees for queries absent from the database $\mathcal{D}$.
1-MIPS Analysis. We begin with top-1 $\mathsf{MIPS}$ of SSG formally as:
Theorem 3.
Consider a vector database $\mathcal{D}\subset\mathbb{R}^{d}$ containing $n$ points, where $n$ is sufficiently large to ensure robust statistical properties. Let $\mathcal{G}$ be a SSG (Fu et al., 2021) defined on $\mathcal{D}$. Assume that the base and query vectors are independently and identically distributed (i.i.d.) and are drawn from the same Gaussian distribution, with each component having zero mean and a variance $\sigma^{2}$. The expected length of the search path $L$ from any randomly selected start node $p$ to a query $q\in\mathbb{R}^{d}$ can be bounded by $\mathbb{E}[L]<c_{0}\frac{\log n+c_{1}d}{\log R+c_{2}d}$, where $R$ is the max-degree of all possible $\mathcal{G}$, independent with $n$ (Fu et al., 2021), and $c_{0},c_{1},c_{2}$ are constants.
Proof Sketch.
Given Theorem 1, 2, along with the monotonic search property of SSG (Fu et al., 2021), Algorithm 1 identifies a greedy search path where each step minimizes the Euclidean distance to $\mu q$ while maximizes the IP distance with respect to $q$, i.e. $\langle r_{i},q\rangle>\langle r_{i-1},q\rangle$. The expected length of the search path can be calculated as:
(4) $\mathbb{E}[L]=\mathbb{E}_{p\in\mathcal{D},q\in\mathbb{R}^{d}}\left[\frac{\langle r_{mip},q\rangle-\langle p,q\rangle}{\mathbb{E}[\langle r_{i},q\rangle-\langle r_{i-1},q\rangle|r_{i-1},q]}\right]$
where $r_{mip}$ is the $\mathsf{MIPS}$ solution of $q$. The outer expectation cannot be simplified directly due to potential dependencies among $p$, $r_{i}$, and $q$. Given that the base and query vectors are finite and drawn from the same distribution, we can enclose them in a hypersphere of diameter $2H$ (Weisstein, 2002), which is independent of $p$, $r_{i}$, and $q$. Since $\langle p,q\rangle<H^{2}$ and $\langle r_{mip},q\rangle<H^{2}$, we can derive:
(5) $\mathbb{E}[L]<\frac{2H^{2}}{\mathbb{E}_{p\in\mathcal{D},q\in\mathbb{R}^{d}}\left[\mathbb{E}\left[\langle r_{i},q\rangle-\langle r_{i-1},q\rangle|r_{i-1},q\right]\right]}$
Although the exact solution to this inequality is intractable, a practical approach involves approximating both the numerator and the denominator. By applying Extreme Value Theory (EVT(Smith, 1990), we can derive a tight upper bound for this approximation, ensuring convergence to the bound as $n$ becomes sufficiently large.
As $H$ is the half-diameter, $H^{2}$ is the maximum squared norm of the dataset $\mathcal{D}$. We define $M_{0}=H^{2}=\max_{1}^{n}\{X_{1}^{2},...,X_{n}^{2}\}$, where $X_{i}$ are element-wise i.i.d. samples from $\mathcal{N}(0,\sigma^{2})$. Consequently, $X^{2}$ follows a chi-square distribution with $d$ degrees of freedom. The Moment Generating Function (MGF(Curtiss, 1942) of $M_{0}$ is given:
(6) $MGF(t)_{X^{2}}=(1-2\sigma^{2}t)^{-\frac{d}{2}},0<t<\frac{1}{2\sigma^{2}}$
From Jensen’s Inequality (McShane, 1937) with $\phi(x)=e^{t_{0}x},t_{0}>0$, we have:
(7) $e^{t_{0}\mathbb{E}[M_{0}]}$ $\leq\mathbb{E}[e^{t_{0}M_{0}}]=\mathbb{E}\max_{i=1}^{n}e^{t_{0}X_{i}^{2}}$
$\leq\sum_{1}^{n}\mathbb{E}\left[e^{t_{0}X_{i}^{2}}\right]=nMGF(t_{0})_{X^{2}}$
(8) $\mathbb{E}[M_{0}]\leq\frac{1}{t_{0}}\left(\log n+\frac{d}{2}\log(1-2\sigma^{2}t_{0})^{-1}\right)$
Here, $t_{0}$ is a hyper-parameter that can be optimized within the range $(0,0.5)$ to enhance the tightness of this bound.
A similar approach can be applied to the random variable $XY$ for approximating the denominator of Eq. 5, where $X$ and $Y$ element-wise i.i.d. sampled from $\mathcal{N}(0,\sigma^{2})$. In this case, $XY$ follows a distribution characterized by a modified Bessel function of the second kind (Bowman, 1958). Then we can have:
(9) $\mathbb{E}[\langle r_{i},q\rangle]\leq\mathbb{E}[\langle r_{i-1},q\rangle]+\frac{1}{t_{1}}\left(\log R+d\log(1-\sigma^{2}t_{1}^{2})^{-1}\right)$
Substituting them into Eq.(5), we have:
(10) $\mathbb{E}[L]$ $<\frac{2H^{2}}{\mathbb{E}_{p\in\mathcal{D},q\in\mathbb{R}^{d}}\left[\mathbb{E}\left[\langle r_{i},q\rangle-\langle r_{i-1},q\rangle|r_{i-1},q\right]\right]}$
$\approx\frac{\frac{2}{t_{0}}\left(\log n+\frac{d}{2}\log(1-2\sigma^{2}t_{0})^{-1}\right)}{\frac{1}{t_{1}}\left(\log R+d\log(1-\sigma^{2}t_{1}^{2})^{-1}\right)}$
where $R$ is the maximum degree of $\mathcal{G}$, independent of $n$ and choices of $p$, $q$, and $r_{i}$(Fu et al., 2021). Thus, the outer expectation can be eliminated. When $t_{0}$ and $t_{1}$ are optimized for the tightest approximation, introducing constants $c_{0}$, $c_{1}$ simplifies the formula to:
(11) $\mathbb{E}[L]<c_{0}\frac{\log n+c_{1}d}{\log R+c_{2}d}$
The complete process of derivation can be found in (Chen et al., 2025). ∎
While Theorem 3 provides an approximation for the expected search path length, the results align well with intuitive insights: the length $L$ increases gradually with $n$ while decreases with graph max-degree $R$. In SSG, the inter-edge angle $\alpha$ governs the graph sparsity (Fu et al., 2021). A smaller $\alpha$ leads to higher $R$, enhancing graph connectivity and thereby reducing inter-node transition costs. Moreover, the overall search complexity for top-1 $\mathsf{MIPS}$ can be bounded by $O\left(c_{0}R\frac{\log n+c_{1}d}{\log R+c_{2}d}\right)$, indicating that the complexity increases more rapidly with $R$ than with $n$, highlighting the critical role of graph pruning for better searching efficiency.
$k$-MIPS Analysis. The preceding theory examines the validity and efficiency of tackling the 1-$\mathsf{MIPS}$ problem on an SSG. Next, We demonstrate that high-confidence $k$-$\mathsf{MIPS}$ solutions can be obtained by exploring the neighborhoods of the 1-$\mathsf{MIPS}$ solution.
Theorem 4.
Given a vector database $\mathcal{D}\subset\mathbb{R}^{d}$ consisting of $n$ points, where $n$ is sufficiently large to ensure robust statistical properties. For the ease of calculation, we assume base and query vectors are element-wise i.i.d. and are drawn from the same Gaussian distribution, parameterized with zero mean and a variance $\sigma^{2}$. Given a tunable parameter $s>0$, we have $\langle r,q\rangle$ is lower bounded by $\left|\lVert p\rVert-s\right|\lVert q\rVert\cos{\left(\theta_{pq}+2\sin^{-1}\frac{s}{2\lVert p\rVert}\right)}$ with probability $Q(s)=\frac{\gamma\left(\frac{d}{2},\frac{s}{2\sigma^{2}}\right)}{\Gamma\left(\frac{d}{2}\right)}$, where $\Gamma(\cdot)$ is the gamma function and $\gamma(\cdot)$ is the lower incomplete gamma function.
Proof.
For any query $q$ and any point $r\in\mathcal{D}$, this proof aims to establish their relationship regarding $q$’s top-1 $\mathsf{MIPS}$ solution $p$, We begin by expand $\langle r,q\rangle$ for $r$ and $q$ as:
(12) $\langle r,q\rangle=\lVert r\rVert\lVert q\rVert\cos{\theta_{rq}}$
, where $\theta_{rq}$ is the angle between vector $r$ and $q$. Under the same norm of $r$, $\cos{\theta_{rq}}$ w.r.t. $p$ is minimized when normalized $r$, $p$, and $q$ fall on the same great circle of the unit sphere. According to spherical trigonometry cosine rule, above equation is bounded as:
(13) $\langle r,q\rangle\geq\lVert r\rVert\lVert q\rVert\cos{(\theta_{rp}+\theta_{pq})}$
With triangle inequality, we further have:
(14) $\langle r,q\rangle\geq\left|\lVert p\rVert-\Delta_{pr}\right|\lVert q\rVert\cos{(\theta_{rp}+\theta_{pq})}$
, where $\Delta_{pr}$ is the $L_{2}$ distance between point $p$ and $r$. We can bound $\Delta_{pr}$ according to the distribution of $\Delta_{pr}$. Specifically, let $Z=\lVert X-Y\rVert^{2}$ denote the distribution of Euclidean distance between $X$ and $Y$, i.i.d. drawn from $\mathcal{D}$. $Z$ follows a gamma distribution parameterized as :
(15) $Z=\lVert X-Y\rVert^{2}\sim Gamma\left(\frac{d}{2},2\sigma^{2}\right)$
The probability that $\mathbb{P}[Z<s]$ for a given threshold $s$ can be derived from the Cumulative Density Function (CDF) as:
(16) $\mathbb{P}[Z<s]=Q(s)=\frac{\gamma\left(\frac{d}{2},\frac{s}{2\sigma^{2}}\right)}{\Gamma\left(\frac{d}{2}\right)}$
With a given $\Delta_{pr}$, the maximum of $\theta_{rp}$ can also be calculated as $\max\theta_{rp}=2\sin^{-1}\frac{s}{2\lVert p\rVert}$. Combining Equation (14) and (16), we derive the lower bound for $\langle r,q\rangle$ as :
(17) $\langle r,q\rangle>\left|\lVert p\rVert-s\right|\lVert q\rVert\cos{\left(\theta_{pq}+2\sin^{-1}\frac{s}{2\lVert p\rVert}\right)}$
, with probability $Q(s)$ in Equation (16). ∎
According to Theorem 4, reducing $s$ drives the convergence of $\langle r,q\rangle$ towards $\langle p,q\rangle$, while also decreasing the estimated number of nodes ($nQ(s)$) that maximize the IP values relative to $q$. Since $s$ represents the upper bound of the Euclidean distance between $r$ and $q$, a judicious choice of $s$ can effectively identify about $nQ(s)$ points in $\mathcal{D}$ as both the top-$k$ $\mathsf{MIPS}$ solutions for $q$ and the range-$s$ nearest neighbors of $p$. In most Euclidean proximity graphs, it is likely that neighbors of neighbors are also neighbors (Prokhorenkova and Shekhovtsov, 2020). Thus, after locating the top-1 $\mathsf{MIPS}$ solution $r_{0}$ for query $q$, Algorithm 1 may require extra iterations to navigate among $r_{0}$’s nearest neighbors for the remaining $k-1$ solutions. Due to the efficient pruning in sparse graphs like SSG, the nearest neighbors of $r_{0}$ are expected to be no more than $O(f(n)n^{1/d}\log n)$ hops away (Fu et al., 2019), where $f(n)$ is a slowly increasing function of $n$. Thus, the search complexity for top-k $\mathsf{MIPS}$ is bounded by $O\left(c_{0}\frac{\log n+c_{1}d}{\log R+c_{2}d}+c_{3}f(n)kRn^{1/d}\log n\right)$, where $c_{3}$ is a scaling constant that integrates the terms linearly.
Figure 2. An illustration of biased solutions for MIPS problems. Spherical Pathways benefit the routing from the inner to the outer ring and the spherical spanning for k answers.
Input: Dataset $\mathcal{D}$, candidate size $L$, maximum degree $R$, minimum angle $\alpha$, refine edge quota $S$, navigation sample number $n$, $k$NNG neighbor number $K$.
Output: $\mathsf{PSP}$ index $G$
1
[…]
3 for each node $i$ in $G$ do
4       $G\leftarrow$ EF$(i,S,\alpha,G)$; $\triangleright$ refer to Alg. 2
5
6N $\leftarrow$ SN$(\mathcal{D},n)$triangleright refer to Alg. 3
7 $G\leftarrow\{G,N\}$
8
9return $G$
Algorithm 2 Graph Index Construction
Remarks. Previous transformation-based methods (Shrivastava and Li, 2014, 2015; Neyshabur and Srebro, 2015; Li et al., 2018; Yan et al., 2018; Zhao et al., 2023a; Zhou et al., 2019) sought to reduce the $\mathsf{MIPS}$ problem to the $\mathsf{NNS}$ problem using non-linear transformations, suffering from strong theoretical assumptions (Zhou et al., 2019), structural distortions (Zhao et al., 2023a; Morozov and Babenko, 2018), and complicate data updates (§2), while failed to provide theoretical guarantees on the reachability of the 1-$\mathsf{MIPS}$ solution. Our Theorems 1 and 2 demonstrate that this reachability can be achieved without space transformations. Theorems 3 and 4 validate the efficiency of solving $\mathsf{MIPS}$ on an SSG, highlighting the crucial role of edge pruning in search performance — a factor which is previously overlooked(Morozov and Babenko, 2018; Liu et al., 2020). Our extensive experiments in §3 substantiate this analysis.
4. Practical Solution
In this section, we introduce a novel $\mathsf{MIPS}$ framework tailored for practical applications. This framework comprises two key components: a new graph-based indexing structure named Proximity Graph with Spherical Pathways (PSP) and a novel search algorithm with Adaptive Early Termination (AET). The PSP index is specifically designed to address the Biased Solution problem prevalent in $\mathsf{MIPS}$ approaches. Concurrently, the AET mechanism effectively mitigates the Excessive Exploration problem, enhancing search efficiency across diverse query distributions.
4.1. Proximity Graph with Spherical Pathway
Base Structure. Advanced theoretical models for nearest neighbor search often exhibit $O(N^{2}\log N)$ indexing time complexity (Fu et al., 2019, 2021), which impedes practical application of our theoretical insights. To mitigate this, we employ an approximate Euclidean proximity graph, NSSG, a practical adaptation of its theoretical counterpart SSG (Fu et al., 2021). This choice is driven by its efficiency and our theoretical analysis in §3. Please refer to SSG paper for details in NSSG indexing algorithm.
Input: Node $n$, new edge quota $S$, angle $\alpha$, NSSG $G$.
Output: A refined graph $G^{\prime}$
1 $E_{n}\leftarrow\emptyset$; $C\leftarrow$ 2-Hop neighbors of n on $G$; $G^{\prime}\leftarrow G$
2
3sort $C$ in descending order of $\langle n,c_{i}\rangle$, $c_{i}\in C$;
4
[…]
6 while $C$ is not empty do
7       $p\leftarrow$ $C[0]$; $C$.remove($C[0]$);
8
9      if $cos\angle nop>\alpha$ then
10             continue;
11      $E_{n}$.add($(n,p)$);
12       if $E_{n}$.size() $\geq S$ then
13             break;
14
15for each edge $e$ in $E_{n}$ do
16       $G^{\prime}$.add($e$) $\triangleright$ deduplicate if repeated
return $G^{\prime}$
Algorithm 3 Edge Refinement (EF)
Input: Dataset $\mathcal{D}$, navigation sample number $m$.
Output: Inverted File $N$
1 $N\leftarrow\emptyset$;$\mathcal{D}^{\prime}\leftarrow$ Sample($\mathcal{D}$) $\triangleright$ cluster on subset $\mathcal{D}^{\prime}$ for big $\mathcal{D}$
2 $C\leftarrow$ Normalized k-means$(\mathcal{D},\mathcal{D}^{\prime},c)$; $\triangleright$ c is cluster number
3 $m_{c}\leftarrow m/c$;
4 for $i\in range(0,c)$ do
5       $S_{m}\leftarrow$ sampling $m_{c}$ nodes from $C[i]$ with Gaussian;
6       $N[i]$.add$(S_{m})$;
7
8
return $N$
Algorithm 4 Spherical Navigation (SN)
Proximity Graph with Spherical Pathways (PSP) addresses the biased solution problem by injecting new edges into the SSG, named as Spherical Pathways (SP), and introducing a lightweight extra structure for search entry management, named as Spherical Navigation (SN). These two components are introduced as follows.
Solving the $\mathsf{MIPS}$ problem often results in biased solutions towards points in regions of large norms, as highlighted in  (Liu et al., 2020). Traditional proximity graphs under the Euclidean metric typically do not inherently address this bias, as their edge selection criteria focus on local neighborhood connectivity, leading to limited sensory regions (Wang et al., 2021b). Empirical observations (Figure 2) reveal that $\mathsf{MIPS}$ routing typically involves initially hopping towards a few answers from inner to outer rings, followed by a broader exploration process for the remaining answers. To enhance this process, we introduce Spherical Pathways, which benefit the $\mathsf{MIPS}$ routing in two aspects (Figure 2): (1) Linking each node to its $\mathsf{MIPS}$ neighbors in the outer ring to accelerate the inner-out routing; (2) Strengthening mutual connectivity within the ”outer ring”, thereby expanding sensory regions crucial for the spanning process in $k$-$\epsilon$$\mathsf{MIPS}$.
Identifying $\mathsf{MIPS}$ neighbors for each node on a large scale is challenging; we address this with a heuristic that involves collecting $k$-hop neighbors and selecting up to $S$ nodes with the largest IP distances as new neighbors. To widen the sensory region, we introduce a smallest-angle constraint inspired by the techniques used in NSSG indexing (Fu et al., 2021). Specifically, any new $\mathsf{MIPS}$ neighbor $m$ of node $n$ (excluding the one with the largest IP distance), the angle $\angle nom\geq\alpha$. Empirically, we find that collecting 2-hop neighbors has significantly enhanced search performance by 8% (§3). This process, termed Edge Refinement (EF), is detailed in Algorithm 3.
Figure 3. An illustration of k-means-based navigation node selection w/ and w/o normalization. Squared with dark edges are randomly selected navigation nodes with large norms.
Spherical Navigation (SN). Navigation nodes (entry nodes of the search algorithm) are pivotal in optimizing search performance on proximity graphs (Fu and Cai, 2016; Malkov and Yashunin, 2018). Given the bias in $\mathsf{MIPS}$ solutions towards large-norm vectors, selecting navigation nodes among them can be advantageous. To prevent over-concentration of navigation nodes, we cluster the base vectors’ normalised projections on a unit sphere using $k$-means, prioritizing angular separation over Euclidean distance (see Figure 3). We then sample $m/c$ navigation nodes from each of the $c$ clusters based on their norm distribution in original space. While the norm distribution follows a long-tail pattern, the Gaussian distribution can serve as an effective approximation in practice. The resulting structure maintaining the selected navigation nodes is encapsulated as an Inverted File (IVF(Jegou et al., 2010). Clusters are keyed by their centers, while navigation nodes are stored in corresponding inverted lists. The SN selection process is outlined in Algorithm 4. Notably, while SN is built on normalised vectors, it only generates the entry points’ IDs. IP metric is used in the original space when searching on the graph, aligning with our theory.
To summarize, building a PSP (detailed in Algorithm 2) involves three steps: (1) Build an NSSG; (2) Apply EF; (3) Apply SN.
4.2. Searching on PSP
Figure 4. An illustration of early termination intuitions. Figure (a) shows the quantity distribution of visited nodes to reach 99% recall@100 on Text2Image1M. Figure (b) presents the changing trends of four selected features during search for a query. The green dotted lines indicate the optimal stop state. After that (yellow curves), the recall@100 stops increasing.
Standard GNNS (Prokhorenkova and Shekhovtsov, 2020) does not fully release the potential of PSP. To optimize its efficiency for $\mathsf{MIPS}$ and ensure effective alignment with $\mathsf{PSP}$, we employ three key adaptations: (1) selecting navigation nodes based on SN; (2) altering the metric as inner product; and (3) integrating adaptive early termination (Algorithm 5).
Entry Points Selection. Unlike the random initialization used in GNNS, we generates navigation nodes from SN. For a query $q$, we first identify the closest cluster in the SN-IVF with cosine similarity, then randomly fills the initial pool from the corresponding inverted list. This is efficient because of the limited number of clusters, ensuring a swift and precise start to the search.
Altering the Metric. Leveraging Theorem 1 and 2, IP metric is enabled for GNNS on the PSP. Meanwhile, the max-heap and min-heap are interchanged adaptively due to the pursuing of answers with larger IP distances instead of smaller $L_{2}$ distances.
Input: Dataset $\mathcal{D}$, $\mathsf{PSP}$ $G$, query $q$, adaptive early termination function $ET(\cdot)$, candidate set size $l_{s}$, result set size $k$, query scale factor $\mu$.
Output: Top $k$ result set $R$.
1 $R\leftarrow\emptyset$; $Q\leftarrow\emptyset$; $C\leftarrow$ closest navigation cluster look-up;
2 $P\leftarrow$ random sample $l_{s}$ nodes from $C$;
3 for each node $p$ in $P$ do
4       $Q$.add($p$,$\langle p,q\rangle$);
[…]
6 while $Q$.size() do
7       $p\leftarrow Q.pop()[0]$; $R.insert((p,\langle p,q\rangle))$
8       $f_{p}\leftarrow$ get_features($p$); $\triangleright$ calculate features for $ET(\cdot)$
9       if $ET(f_{p})$ then
10            break; $\triangleright$ $ET(\cdot)$ is true for stop
11      if visited($p$) then
12            continue;
13      $N_{p}\leftarrow$ neighbors of $p$ in $G$; $Q\leftarrow$ batch_insert($Q,N_{p}$);
14       $Q.resize(l_{s})$;$R.resize(k)$; $\triangleright$ delete small-IP value nodes
15
16
return $R$
Algorithm 5 Search On PSP
Figure 5. A showcase of a trained decision tree. The crosses denote pruned branches. The yellow nodes predict ”stop”. The conditional expressions are learnt decision logic based on selected features. ”Pos/Neg” denotes the number of positive/negative samples fall in leaves, indicating stop tendency.
Adaptive Early Termination. Early termination is crucial to search efficiency (Liu et al., 2021; Li et al., 2020; Zhang et al., 2023b). The exploration process is highly sensitive to the query distribution, yet standard GNNS is unaware of different query distributions and the accuracy of the result set. As shown in Figure 4(a), the distribution of the number of nodes visited exhibits a long tail, with most queries reaching 99% recall with a small exploration depth. Instead of a uniform configuration for search parameters, adaptively stopping the search earlier for simpler queries can significantly enhance the amortized processing time and reduce excessive explorations. To equip IP-GNNS with ”self-awareness”, our method integrates a decision function to monitor key features indicative of the optimal stopping point, namely Adaptive Early Termination (AET), incorporates the following:
Feature Selection. Features for the decision model are chosen based on the following principles. (1) Relevance: Features must be closely related to the IP metric, the query, and the tendency toward termination. (2) Distinguishability: Features must clearly distinguish between states before and after the termination point. (3) Simplicity: The computational cost of the features should be minimized to maintain efficiency. Theorem 4 indicates the vector norms, the IP values, and exploring neighborhood of top-1 $\mathsf{MIPS}$ solution are key aspects in feature engineering. Notably, the update frequency of the top-k candidate pool serves as a good indicator of the neighborhood exploration. After extensive and rigorous experiments (Figure 4(b)), we finalized the feature set, detailed in Table 1.
Table 1. Features used in adaptive early termination. EMA means exponential moving average.
Feature Description
F1: IP value EMA of (IP distance to $q$)
F2: Norm/Min Norm EMA of (norm / historical min norm)
F3: IP/Max IP EMA of (ip / historical max ip)
F4: Update Times Top k update frequency (using EMA)
Decision Model. Utilizing the identified four highly discriminative features, a simple decision tree (DT) is employed, trained on labeled data points collected via searching on $\mathsf{PSP}$. The data is generated through the following steps: (1) Divide the base vectors into new base and new query sets; (2) Perform searches on the new queries using the base indexed by $\mathsf{PSP}$. For each query, record the node ID (boundary node) when the recall of the result set ceases to increase; (3) Randomly sample nodes before the boundary node (labeled as 1) and after the boundary node (labeled as 0) to form the training data. (4) Train the DT on the collected training data.
Function Generation. To facilitate efficient implementation of the AET decision function, it is essential to convert the DT into C++ code. We propose a structured approach to simplify the DT while maintaining its effectiveness and efficiency: (1) Limit the height of the DT to the number of features to prevent over-fitting. (2) Adjust the prediction mechanism of the leaf nodes (Figure 5). A leaf node predicts ”stop” only when the ratio #$\frac{\#Negative\,Samples}{\#Positive\,Samples}>\theta$. This approach prioritizes higher confidence in stopping the search, reducing the risk of prematurely terminating. Users can adjust $\theta$ to control the aggressiveness of the AET function. (3) Remove subtrees whose leaf nodes share the same predictions to simplify the tree. (4) Translate the resulting DT into ’if-else’ clauses in C++ from the tree structure. Figure 5 is an example of this process. The final interpreted clause is ”if (F4¡0.5 & F3 ¡ 0.6) stop;else continue;”
Remarks. Both candidate size $l_{s}$ and early-stop features affect the search recall by pruning the search space. Considering the monotonicity of these indicators w.r.t. search recall, one can tune the candidate size or the decision boundaries of the leaf nodes manually to get different recall, free of retraining the DT for each recall level.
4.3. Complexity Analysis
The Indexing algorithm consists of two main stages: (1) NSSG indexing and (2) the EF and SN. Let $n$ denote the dataset size, $r$ the max degree of the graph, $d$ the dimension, $c$ the number of clusters in SN, and $m$ the number of navigation nodes in SN. The time complexity of NSSG indexing has an empirical complexity of $O(dn\log n)$ (Fu et al., 2021). For each node, the EF phase involves IP calculation ($O(ndr^{2})$), sorting ($O(nr^{2}logr)$) and new neighbor selection ($O(nr^{2})$), cumulatively yielding $O(nr^{2}(d+logr))$. The SN phase, dominated by the $k$-means cluster assignment, scales as $O(ncd)$. Absorbing smaller constants into $c_{4}$, the overall indexing complexity of $\mathsf{PSP}$ is $O(c_{4}n\log n+n)$, dominated by $O(n\log n)$ term.
The Search complexity for generally favored top-k $\mathsf{MIPS}$ is approximated by $O(c_{0}\frac{\log n+c_{1}d}{\log R+c_{2}d}+c_{3}f(n)kRn^{1/d}\log n)$ in §3. It is mainly dominated by the $O(kRn^{1/d}\log n)$ term. Empirical evaluations (detailed in §5) aligns with our estimation.
Worst Cases. The worst case will invalidate any acceleration techniques for approximate methods, e.g., points on a straight line. In such cases, the search complexity on $\mathsf{PSP}$ will downgrade to near $O(n)$. The indexing complexity of $\mathsf{PSP}$ will downgrade to $O(n^{2})$.
5. Experimental Evaluation
Table 2. Dataset statistics. Dim indicates vector dimension.
Dataset Base Size Dim Query Size Modality
MNIST (mni, 1998) 60,000 784 10,000 Image
DBpedia100K (Oliva and Torralba, 2001) 100,000 3072 1,000 Text
DBpedia1M (Oliva and Torralba, 2001) 1,000,000 1536 1,000 Text
Music100 (Morozov and Babenko, 2018; Hu et al., 2008) 1,000,000 100 10,000 Audio
Text2Image1M (tex, 2021) 1,000,000 200 100,000 Multi
Text2Image10M (tex, 2021) 10,000,000 200 100,000 Multi
Laion10M (Radford et al., 2021b) 12,244,692 512 1,000 Multi
Commerce100M 100,279,529 48 64,111 E-commerce
We conduct extensive experiments on real-world datasets to validate the theoretical contributions of our proposed framework, as well as its efficiency, scalability, and practical applicability for the $\mathsf{MIPS}$ task. Our analyses are guided by the following questions:
  • Q1: How does the proposed theory align $\mathsf{MIPS}$ with $\mathsf{NNS}$?
  • Q2: How does $\mathsf{PSP}$ perform compared to established baselines across various modalities, dimensionalities, and cardinalities?
  • Q3: Ablation study of the search efficiency of $\mathsf{PSP}$.
  • Q4: How is the viability of $\mathsf{PSP}$ in large-scale applications?
5.1. Experimental Setup
Datasets. All empirical evaluations were conducted on eight real-world datasets, covering diverse modalities, cardinalities and dimensionalities (Table 2): MNIST (mni, 1998) and Music100 (Morozov and Babenko, 2018; Hu et al., 2008) datasets are the established benchmarks for the $\mathsf{MIPS}$ problem. DBpedia (Oliva and Torralba, 2001) is extracted by OpenAI text-embedding-3-large model. Text2Image (tex, 2021) is a cross-modality dataset where image embeddings (base) are extracted from Se-ResNext-101 and text embeddings (query) from a DSSM Model. Laion10M (Radford et al., 2021b) is a cross-modal retrieval dataset. The embeddings are generated by the CLIP-ViTB/32 model. Commerce100M, generated from real traffic logs of a large-scale e-commerce platform. The embeddings are via an advanced method (Fu et al., 2024). The dimension is set to 48 due to the application demands in industrial scenarios, while 16 to 64 are common choices in this field (Zhang et al., 2020; Wang et al., 2021a), discussed in (Chen et al., 2025).
Competitors were selected from advanced in-memory methods, encompassing (1) ip-NSW (Morozov and Babenko, 2018): A graph based method using inner-product navigable small world graph. (2) ip-NSW+ (Liu et al., 2020): An enhancement of ip-NSW that introduces an additional angular proximity graph. (3) Möbius-Graph (Zhou et al., 2019): A graph based method that reduces the $\mathsf{MIPS}$ problem to an $\mathsf{NNS}$ problem using Möbius transformation. (4) NAPG (Tan et al., 2021): the state-of-the-art graph based method. (5) Fargo (Zhao et al., 2023a): The latest state-of-the-art LSH based method with theoretical guarantees. (6) ScaNN (Guo et al., 2020): The latest state-of-the-art quantization method, an optimised version based on SOAR (Sun et al., 2024).
Implementation. All baselines except for ScaNN are written in C++. The ScaNN library is called by Python bindings yet written in C++. The experiments are executed on a CentOS machine with 128G RAM on Intel(R) Xeon(R) CPU E5-2650 v4 @ 2.20GHz CPU. We use OpenMP for parallel index construction, utilizing 48 threads for all methods. For query execution, we turn off additional optimizations.
Table 3. The average overlap ratio of NNS-MIPS search paths and recall@100 with varying $\mu$, on MNIST.
Metric $\mu=0.1$ $\mu=1$ $\mu=10$ $\mu=100$ $\mu=1200$
Overlap ratio 7.4% 15.3% 72.4% 99.5% 100%
Recall@100 0.0 0.08 0.83 1.0 1.0
Parameter Settings. $\mathsf{PSP}$ required tuning several hyper-parameters: number of neighbors in $k$NN graph ($K$), NSSG candidate size ($L$), maximum out-degree ($R$), minimal angle between edges ($\alpha$), and the number of nodes added in EF ($S$). By default, we set $K=400$, $L=800$, $R=40$, $\alpha=60^{\circ}$, and $S=5$ for all datasets. For a fair comparison, we also employ grid search to optimize the parameters of baseline methods over all datasets.
Evaluation Metrics. For search performance comparison, we measure effectiveness and efficiency using two popular metrics: (1) Recall vs. Queries Per Second (QPS), denoting the number of queries that an algorithm can process per second at each $recall@100$ level; and (2) Recall vs. Computations, indicating the number of distance computations performed by the search algorithm. Let $R_{t}^{\prime}$ denote the set of $k$ vectors returned by the algorithm, and $R_{t}$ represent the ground truth, the recall@k is formally defined as $recall@k=\frac{|R_{t}\cap R_{t}^{\prime}|}{|R_{t}|}=\frac{|R_{t}\cap R_{t}^{\prime}|}{k}$ Additionally, we consider construction time and index size to evaluate the scalability in practice.
Experimental Results Exp-1: Theory Verification (Q1). To explore the $\mathsf{MIPS}$-$\mathsf{NNS}$ equivalence with respect to $\mu$, we conduct GNNS on an ideal $\mathsf{PSP}$ index, obtained by setting $L$ to the size of the base vectors and $R$ to infinity in Algorithm 2. This experiment is conducted on MNIST, and we also evaluate Theorem 4 on synthetic datasets sampled from a standard normal distribution, examining the relationship between top-k $\mathsf{MIPS}$ solutions and the top-1 solution’s Euclidean neighbors.
Duality of NNS and MIPS on Proximity Graph. We assessed the average overlap ratio of search paths for $\mathsf{NNS}$ and $\mathsf{MIPS}$, and recall@100 for $\mathsf{MIPS}$, across all queries on MNIST under different $\mu$ (Table 3). The result indicates that a sufficiently large $\mu$ can effectively unify the $\mathsf{MIPS}$ and $\mathsf{NNS}$ paths for all queries. Even sub-optimal $\mu$ yields reasonably good candidates due to highly overlapping paths. Case study in (Chen et al., 2025) visualizes this. Additionally, we evaluate the overlap ratio between the top-k $\mathsf{MIPS}$ solutions and the neighborhood of the top-1 $\mathsf{MIPS}$ solution. The results in (Chen et al., 2025) confirm that top-k $\mathsf{MIPS}$ solutions are likely to be distributed in the Euclidean neighborhood of the top-1 $\mathsf{MIPS}$ solution.
Exp-2: Performance Assessments (Q2). The results are presented in Figure LABEL:fig:exp-qps and LABEL:fig:memory. The absence of methods in the figures indicates either high processing time or low recall.
Query Processing Performance. Figure LABEL:fig:exp-qps presents the query processing performance for different methods. The top rows depict QPS at varying recall levels, where upper right is better. The bottom rows show the number of distance computations required, with lower right being better. Log scales are applied for wide value ranges.
PSP consistently outperforms all baselines across all datasets, achieving up to 37% speedup over Möbius-Graph ($2^{nd}$ best) on Laion10M. This demonstrates $\mathsf{PSP}$’s superior efficiency with high-dimensional and large-scale data, leveraging its strong theoretical foundations and ensuring robust connectivity. In contrast, other graph-based methods suffer from connectivity issues, and non-graph methods’ inefficiency mainly owes to inefficient indexing large-norm points.
PSP-, representing an unoptimized NSSG, still surpasses other baselines across datasets by an average of 20%-25%, except on Music100. This aligns with our analysis that MIPS can be executed efficiently without transformation. Further optimizations, like EF, SN, and AET, improve performance by about 10%-15%.
Figure 6. Ablation study results across five different scale datasets. We measure the benefits of three optimization techniques on query execution.
Others: (1) Fargo is inferior to those of graph-based algorithms, notably achieving only 90% recall on Commerce100M. Its reliance on LSH and susceptibility to transformation-induced distortion limits its general performance. (2) ScaNN excels on DBpedia100K but struggles elsewhere, highlighting the sensitivity of quantization methods. For datasets not well-suited to quantization, such as large-scale or IID distributions, performance declines (Li et al., 2019), as seen that its recall is limited on Commerce100M. (3) ip-NSW and ip-NSW+ lag behind $\mathsf{PSP}$ due to high graph density and precision limitations. Although ip-NSW+ partially addresses precision issues, its performance can be inconsistent, especially on datasets like Text2Image1M, where the additional angular-graphs hurt performance. (4) Möbius-Graph shows strong performance on Commerce100M but encounters precision bottlenecks on Music100 and DBpedia1M due to violations of transformation-induced assumptions (§2). (5) NAPG performs similarly to or worse than ip-NSW due to inefficient graph sparsification techniques and lack of theoretical support in their heuristics.
Indexing Performance. Figure  LABEL:fig:memory shows the index memory footprint and indexing time for various methods. PSP achieves a balanced memory footprint and indexing time across all datasets. On Commerce100M, PSP requires only 12 GB and 13 hours for indexing, making it feasible for daily updates—essential for adapting to evolving business metrics. In contrast, ip-NSW+ takes over a day for indexing. While Fargo is the fastest and least memory-intensive, it suffers from poor query performance. Indices of ScaNN, ip-NSW, ip-NSW+, Möbius-Graph, and NAPG consume 1.5 to 2$\times$ the memory of PSP, reducing their practicality. Graph-based methods have an index size about 3$\times$ larger than PSP, excluding data, due to high redundancy in their edges. ScaNN also requires extensive memory for large tree structures to improve search precision. Table 4 illustrates the graph features of PSP across different datasets. Utilizing its theoretical foundations, PSP exhibits a low average out-degree, a moderate clustering coefficient, and strong connectivity.
Table 4. Graph feature statistics of PSP on different datasets.
Datasets Index degree Cluster Coeff Shortest path
Avg. Std. Avg. Std. Avg. Std.
MNIST 29 1.7 0.09 0.002 4.4 0.72
DBpedia100K 35 0.02 0.03 0.004 3.7 0.51
DBpedia1M 35 0.26 0.03 0.006 4.8 0.53
Music100 21 3.7 0.10 0.008 5.3 0.78
Text2image1M 39 1.5 0.05 0.001 5.9 0.9
Text2image10M 39 1.6 0.04 0.009 7.6 1.3
Laion10M 20 6.32 0.11 0.03 7.1 1.6
Commerce100M 30 5.2 0.04 0.009 8.9 1.49
Table 5. The querying and indexing performance over various data scales on Text2image10M at 99% recall.
Dataset Scale Query Time Indexing Time Index Size
Text2image100K 0.13 (ms) 62 (s) 15 (MB)
Text2image1M 0.46 (ms) 498 (s) 148 (MB)
Text2image10M 0.95 (ms) 5672 (s) 1520 (MB)
Exp-2 Summary. PSP’s superiority is attributed to : strong graph connectivity, no transform in space, efficient edge pruning, and tailored adaptations for MIPS. Superior search performance and efficient indexing makes PSP well-suited for large-scale applications.
Exp-3: Ablation Study (Q3). We validates the impact of key components of $\mathsf{PSP}$ across five datasets. Results are recorded at recall@100 level of 99% (Figure 6). Key findings include:
Effect of Spherical Navigation. SN improves average performance by 6.2% across five datasets by starting searches in high-norm areas, reducing redundant computations. On Music100, it provides a 15% speedup, suggesting a highly biased distribution of $\mathsf{MIPS}$ solutions.
Effect of Edge Refinement. EF adds a 3.7% average improvement, and 8% on Music100, by enhancing connectivity and increasing graph degree slightly, especially towards high-norm regions. The ”Spherical Highways” introduced by EF serve as shortcuts, improving navigation efficiency. The effects of EF and SN may overlap, reducing additional benefits when combined.
Effect of Adaptive Early Termination. AET yields an average improvement of 5.1%, mitigating redundant paths once the optimal set is found. It is particularly effective for queries following long-tailed distributions, with improvements of 10.6% on Text2Image1M and 8.3% on DBpedia100K. Further speedups are achievable via smaller $\theta$ setting (§4.2) with minor acceptable recall loss.
Figure 7. Query time versus data scale on Text2image at 99% recall for top-1 and top-k $\mathsf{MIPS}$. For top-1 queries, both the $x$ and $y$ axes use a logarithmic scale, resulting in an almost straight line, indicating a growth rate close to $O(\log n)$.
Exp-4: Application Viability (Q4) is determined by several aspects: the ease of hyper-parameter tuning, the scalability with dataset cardinality, the scalability with demanded answer number (k), the ease of tuning recall@k level, the consistency of query variance, and the flexibility to support diverse search scenarios. We find: (1) Extensive tests on Text2Image confirm the theoretical complexity estimates from §3. Results (Figure 7) show $O(\log n)$ scalability for top-1 $\mathsf{MIPS}$ and near $O(\log n)$ scalability for top-k $\mathsf{MIPS}$, with manageable indexing times and approximately linear index size growth (Table 5). PSP achieves 99% recall with a query time of just 0.9 ms on a 10 million scale dataset, demonstrating strong scalability with a small constant factor. (2) The performance of $\mathsf{PSP}$ is mainly determined by $R$ and quality of the base NSSG. Generally, higher quality of NSSG leads to better performance. Optimal $R$ exists yet remains consistent across all scales, allowing for efficient tuning on smaller subsets. (3) The query process time grows near sub-linearly with $k$, allowing for efficient search with large return quantity. (4) The query performance demonstrates strong robustness across different datasets (Table 6). The variance is small over random runs. (5) PSP can seamlessly support cosine similarity search (a special case of MIPS), allowing it to be widely adopted by various search vendors.
Table 6. Mean and standard deviation of the query execution time (ms) over various data scale at 95% recall.
Metric Music100 Text2image10M Commerce100M
time $0.095\pm 0.02$ $0.543\pm 0.09$ $2.631\pm 0.35$
6. Discussion
Strengths of New Dataset. The new Commerce100M dataset is derived from large-scale real online traffic data collected from the Shopee e-commercial search. The vectors are generated with Resflow (Fu et al., 2024), advanced deep learning method in recommender and search. This dataset is unique for several reasons: (1) Commerce100M’s intrinsic dimension closely matches its actual dimension, leading to a distinct spatial topology compared to other ”high-dimensional” datasets. (2) The queries represent users while the base vector represents groceries, a special type of cross-domain. (3) The performance on Commerce100M aligns perfectly with that on Shopee APP, rendering a real-world arena for this literature.
Limitations. (1) This study may not fully capture variability across extremely skewed norm distributions. (2) Our theory does not specifically prune redundant edges under MIPS setting. (3) This work has not yet explored incremental indexing strategies.
Future Works. Building on the identified limitations, future research should focus on three key aspects: (1) Ascertain the adaptability and robustness of our methods across a broader array of heterogeneous datasets. (2) Consider developing acceleration techniques and theory specifically tailored to the MIPS problem. (3) Explore incremental indexing that can adapt real-time updates.
7. Other Related Works
Inner Product (IP) is ubiquitous in metric learning, classification, clustering, knowledge graphs, recommender system, large language model, and information retrieval (Yu et al., 2014; Fujiwara et al., 2023; Xu et al., 2020; Huang et al., 2020; Grbovic and Cheng, 2018; Pan et al., 2023; Zhao et al., 2023b; Wang et al., 2021c; Guo et al., 2022; Wang et al., 2024) MIPS methods can be categorized by their indexing strategies into LSH, tree, quantization, and graph based approaches:
LSH-based methods. Adapting LSH to the IP metric includes L2 (Shrivastava and Li, 2014), Correlation (Shrivastava and Li, 2015), and the popular XBOX (Bachrach et al., 2014), which introduces certain data distortion. Among the works based on XBOX (Neyshabur and Srebro, 2015; Keivani et al., 2018; Pham, 2021; Yan et al., 2018; Huang et al., 2018; Zhao et al., 2023a), Fargo (Zhao et al., 2023a) achieves state-of-the-art.
Tree-based methods. Early tree-based MIPS research (Ram and Gray, 2012; Koenigstein et al., 2012) struggled with high dimensionality. ProMIPS (Song et al., 2021) tackles this issue by projecting the data into lower-dimensional spaces but it suffers from severe information loss. LRUS-CoverTree (Ma et al., 2024) was designed to mitigate this, while it struggles with solving negative IP values.
Quantization-based methods target to accelerate distance computations with certain approximations. ScanNN (Guo et al., 2020) combine the ”VQ-PQ” framework with an anisotropic quantization loss, presenting a cutting-edge quantization-based solution. SOAR (Sun et al., 2024) uses an orthogonality-amplified residual loss to enhance representations and reaches state-of-the-art and is part of ScaNN library.
Graph-based methods. Proven effective for NNS (Malkov and Yashunin, 2018; Fu et al., 2019, 2021; Lu et al., 2021; Feng et al., 2023; Azizi et al., 2023), graph-based techniques have been adapted for MIPS. ip-NSW (Morozov and Babenko, 2018) replaces Euclidean metric in NSW (Malkov et al., 2014) by IP, and ip-NSW+ (Liu et al., 2020) further improves it by adding an angular proximity graph. Möbius-Graph (Zhou et al., 2019) introduces Möbius transformation, albeit with assumptions. Other algorithms  (Tan et al., 2019, 2021) introduce heuristic edge selection strategies to accelerate the search.
Other Acceleration Techniques.  (Li et al., 2018) utilizes machine learning to optimize the LSH indices. EI-LSH (Liu et al., 2021) enhances LSH-based methods with early termination techniques. Similar methodologies have also been explored in graph-based methods (Li et al., 2020) for NNS.
8. Conclusion
In this paper, we align Maximum Inner Product Search (MIPS) with Nearest Neighbor Search (NNS) through robust theoretical foundations, presenting the first analytical framework for $\mathsf{MIPS}$ efficiency on graph indices. To address the inherent challenges of biased solutions and excessive explorations in graph-based $\mathsf{MIPS}$, we propose a practical framework, featuring a novel graph index, $\mathsf{PSP}$, and a novel lightweight adaptive early termination mechanism, AET. Extensive experiments validate the theoretical soundness, efficiency, scalability, and practical applicability of our approaches. Notably, our method has been deployed in the Shopee search engine, demonstrating superior performance. Additionally, we contribute the Commerce100M dataset to the public community, addressing the lack of e-commerce data in this domain.
Acknowledgements.
This work was supported in part by the NSFC under Grants No. (62025206 and U23A20296), Zhejiang Province’s “Lingyan” R&D Project under Grant No. 2024C01259, CCF-Aliyun2024004, Ningbo Yongjiang Talent Introduction Programme (2022A-237-G). Cong Fu is the main project advisor. Xiangyu Ke is the corresponding author of this work.
Appendix
Appendix A Complete Proof of THEOREM 2
Given a proximity graph $G=(V,E)$ and a query $q\in\mathbb{R}^{d}$, when using the standard Graph Nearest Neighbor Search ($\mathsf{GNNS}$(Prokhorenkova and Shekhovtsov, 2020) algorithm to decide the $\mathsf{MIPS}$ solution for $q$, there exists a scalar $\bar{\mu}$ such that for $\forall\mu>\max(\bar{\mu},0)$, we can identify a node $p^{*}\in\mathcal{N}_{o}$ that satisfies: $p^{*}\in\left\{\operatorname*{arg\,max}_{p\in\mathcal{N}_{o}}\langle p,q\rangle\right\}\cap\left\{\operatorname*{arg\,min}_{p\in\mathcal{N}_{o}}\delta(p,q^{\prime})\right\}$. Here, $o$ can be any node on the search path of $\mathsf{GNNS}$ regarding the query $q^{\prime}=\mu q$, and $\mathcal{N}_{o}=\left\{p|(p,o)\in E\right\}$ denotes $o$’s neighbor nodes.
Proof.
The search path of the Algorithm 1 on a proximity graph typically expands by selecting a node from the neighbors of the current node that minimizes the Euclidean distance to $q$ for $\mathsf{NNS}$, yet maximizes the inner product for $\mathsf{MIPS}$ (more details can be found in  (Wang et al., 2021b)). To align the search paths under both metrics, we must unify the node-selection behavior along the path.
This challenge is addressed by localizing the problem: solving for the optimal $p^{*}$ among the neighbors at each greedy step, akin to the conditions established in Theorem 1. Let $l$ be the number of steps in the path. Define $U=\left\{\bar{\mu}_{i}|1\leq i\leq l\right\}$, the set of $\bar{\mu}$ solving all sub-problems. We have $\forall\mu>\sup\left(U\cup\{0\}\right)$, Algorithm 1 selects the same node under both IP and Euclidean metrics at each step.
In the special case where multiple neighbors qualify as local $\mathsf{MIPS}$ solutions for $\mu q$, this ambiguity can be resolved by a simple adaptation to the standard Algorithm 1 protocol: always select the nearest neighbor of $\mu q$ with smallest ID among the qualified nodes.
In summary, there exists a scalar boundary $\sup\left(U\cup\{0\}\right)$ which forces the Algorithm 1 under both IP and Euclidean metrics to generate the same search path targeting the scaled query $\mu q$. ∎
Appendix B Complete Proof of THEOREM 3
Consider a vector database $\mathcal{D}\subset\mathbb{R}^{d}$ containing $n$ points, where $n$ is sufficiently large to ensure robust statistical properties. Let $\mathcal{G}$ be a SSG (Fu et al., 2021) defined on $\mathcal{D}$. Assume that the base and query vectors are independently and identically distributed (i.i.d.) and are drawn from the same Gaussian distribution, with each component having zero mean and a variance $\sigma^{2}$. The expected length of the search path $L$ from any randomly selected start node $p$ to a query $q\in\mathbb{R}^{d}$ can be bounded by $\mathbb{E}[L]<c_{0}\frac{\log n+c_{1}d}{\log R+c_{2}d}$, where $R$ is the max-degree of all possible $\mathcal{G}$, independent with $n$ (Fu et al., 2021), and $c_{0},c_{1},c_{2}$ are constants.
Proof.
Given Theorem 1, 2, along with the monotonic search property of SSG (Fu et al., 2021), Algorithm 1 identifies a greedy search path where each step minimizes the Euclidean distance to $\mu q$ while maximizes the IP distance with respect to $q$, i.e. $\langle r_{i},q\rangle>\langle r_{i-1},q\rangle$. The expected length of the search path can be calculated as:
$\mathbb{E}[L]=\mathbb{E}_{p\in\mathcal{D},q\in\mathbb{R}^{d}}\left[\frac{\langle r_{mip},q\rangle-\langle p,q\rangle}{\mathbb{E}[\langle r_{i},q\rangle-\langle r_{i-1},q\rangle|r_{i-1},q]}\right]$
where $r_{mip}$ is the $\mathsf{MIPS}$ solution of $q$. The outer expectation cannot be simplified directly due to potential dependencies among $p$, $r_{i}$, and $q$. Given that the base and query vectors are finite and drawn from the same distribution, we can enclose them in a hypersphere of diameter $2H$ (Weisstein, 2002), which is independent of $p$, $r_{i}$, and $q$. Since $\langle p,q\rangle<H^{2}$ and $\langle r_{mip},q\rangle<H^{2}$, we can derive:
$\mathbb{E}[L]<\frac{2H^{2}}{\mathbb{E}_{p\in\mathcal{D},q\in\mathbb{R}^{d}}\left[\mathbb{E}\left[\langle r_{i},q\rangle-\langle r_{i-1},q\rangle|r_{i-1},q\right]\right]}$
Although the exact solution to this inequality is intractable, a practical approach involves approximating both the numerator and the denominator. By applying Extreme Value Theory (EVT(Smith, 1990), we can derive a tight upper bound for this approximation, ensuring convergence to the bound as $n$ becomes sufficiently large.
As $H$ is the half-diameter, $H^{2}$ is the maximum squared norm of the dataset $\mathcal{D}$. We define $M_{0}=H^{2}=\max_{1}^{n}\{X_{1}^{2},...,X_{n}^{2}\}$, where $X_{i}$ are element-wise i.i.d. samples from $\mathcal{N}(0,\sigma^{2})$. Consequently, $X^{2}$ follows a chi-square distribution with $d$ degrees of freedom. The Moment Generating Function (MGF(Curtiss, 1942) of $M_{0}$ is given:
$MGF(t)_{X^{2}}=(1-2\sigma^{2}t)^{-\frac{d}{2}},0<t<\frac{1}{2\sigma^{2}}$
From Jensen’s Inequality (McShane, 1937) with $\phi(x)=e^{t_{0}x},t_{0}>0$, we have:
$e^{t_{0}\mathbb{E}[M_{0}]}$ $\leq\mathbb{E}[e^{t_{0}M_{0}}]=\mathbb{E}\max_{i=1}^{n}e^{t_{0}X_{i}^{2}}$
$\leq\sum_{1}^{n}\mathbb{E}\left[e^{t_{0}X_{i}^{2}}\right]=nMGF(t_{0})_{X^{2}}$
$\mathbb{E}[M_{0}]\leq\frac{1}{t_{0}}\left(\log n+\frac{d}{2}\log(1-2\sigma^{2}t_{0})^{-1}\right)$
Here, $t_{0}$ is a hyper-parameter that can be optimized within the range $(0,0.5)$ to enhance the tightness of this bound. When $n$ is sufficiently large, this upper bound serves as a good approximation of $\mathbb{E}[M_{0}]$.
A similar approach can be applied to the random variable $XY$ for approximating the denominator of Eq. 5, where $X$ and $Y$ element-wise i.i.d. sampled from $\mathcal{N}(0,\sigma^{2})$. In this case, $XY$ follows a distribution characterized by a modified Bessel function of the second kind (Bowman, 1958). The MGF for $XY$ is given by:
$MGF(t)_{XY}=\left(\sqrt{1-\sigma^{2}t_{1}^{2}}\right)^{-d}$
Define $M_{1}=\max_{0}^{R}\{X_{1},..,X_{R}\}$, where $R$ is the maximum degree of SSG $\mathcal{G}$. Applying Jensen’s Inequality again, we obtain:
$e^{t_{1}\mathbb{E}[M_{1}]}$ $\leq E[e^{t_{1}M_{1}}]$
$\mathbb{E}[M_{1}]$ $\leq\frac{1}{t_{1}}\left(\log R+d\log(1-\sigma^{2}t_{1}^{2})^{-1}\right)$
where $10<\sigma^{2}t_{1}^{2}<$ 1. This derived upper bound assumes a distribution centered at the origin. By applying Re-scaling (shifting), we can derive the bound for the case of neighbors of $r_{i-1}$ (points centered at $r_{i-1}$ in a Euclidean proximity graph) as follows:
$\mathbb{E}[\langle r_{i},q\rangle]$ $=\mathbb{E}\left[\max_{m=1}^{R}\{\langle r_{m},q\rangle|r_{m}\in N(r_{i-1})\}\right]$
$\leq\mathbb{E}[\langle r_{i-1},q\rangle]+\frac{1}{t_{1}}\left(\log R+d\log(1-\sigma^{2}t_{1}^{2})^{-1}\right)$
Substituting them into Eq.(5), we have:
$\mathbb{E}[L]$ $<\frac{2H^{2}}{\mathbb{E}_{p\in\mathcal{D},q\in\mathbb{R}^{d}}\left[\mathbb{E}\left[\langle r_{i},q\rangle-\langle r_{i-1},q\rangle|r_{i-1},q\right]\right]}$
$\approx\frac{\frac{2}{t_{0}}\left(\log n+\frac{d}{2}\log(1-2\sigma^{2}t_{0})^{-1}\right)}{\frac{1}{t_{1}}\left(\log R+d\log(1-\sigma^{2}t_{1}^{2})^{-1}\right)}$
where $R$ is the maximum degree of $\mathcal{G}$, independent of $n$ and choices of $p$, $q$, and $r_{i}$(Fu et al., 2021). Thus, the outer expectation can be eliminated. When $t_{0}$ and $t_{1}$ are optimized for the tightest approximation, introducing constants $c_{0}$, $c_{1}$ simplifies the formula to:
$\mathbb{E}[L]<c_{0}\frac{\log n+c_{1}d}{\log R+c_{2}d}$
Appendix C Details of Commerce100M dataset.
Commerce100M is derived from real traffic logs of a large-scale e-commerce platform, with embeddings generated by an advanced, pre-trained personalized deep retrieval model. The embeddings have a dimensionality of 48, aligning with industrial norms and resource constraints, where vector dimensions typically range from 16 to 64 due to the vast number of candidates and memory limitations (Zhang et al., 2020; Wang et al., 2021a). In this setup, storing 100 million 48-dimensional vectors consumes about 20GB of memory, leaving 30GB for indices under a 50GB per-shard memory allocation for MIPS tasks. Additionally, these vectors are utilized for ranking items among retrieved candidates, with the ranking engine capping the dimension at 64 for efficient similarity ranking.
Query vectors represent joint user and search keyword profiles, while base vectors encompass over 100 million popular grocery items. The query vectors are categorized into three types. Each vector is composed of three segments, each containing 16 dimensions, used in different combinations for various similarity calculations reflecting user behaviors:
  • When using part1 (16 dimensions) of the query and part1 (16 dimensions) of the item for similarity calculation, it indicates the user’s click tendency.
  • When using the concatenation of part1 and part2 of the query with the concatenation of part1 and part2 of the item for similarity calculation (32 dimensional vector), it indicates the user’s add-to-cart tendency.
  • When using the concatenation of part1, part2, and part3 of the query with the concatenation of part1, part2, and part3 of the item for similarity calculation (48 dimensional vector), it indicates the user’s purchase tendency.
Moreover, each query contains a sequence of items with which the user has actually interacted and their corresponding labels. A query may correspond to one or multiple interacted items. Different labels represent different levels of interaction:
  • label=3: The user only clicked.
  • label=4: The user clicked and added to the cart.
  • label=5: The user clicked, added to the cart, and made a purchase.
Insights from deploying PSP: Shopee’s e-commerce retrieval application is driven by specific real-world business needs. (i) On the Shopee platform, a large number of products are batch-listed and delisted within a brief time frame. (ii) Considering the limitations of online service resources and the requirements for multi-replica disaster recovery and traffic distribution, we must constrain the index update time, the memory utilized for index updates, and the memory required for index serving. Many existing methods either fail to meet the retrieval efficiency requirement (e.g., ScaNN), or consume excessive index resources. In contrast, PSP perfectly matchs all our requirements.
Appendix D Additional results for Fargo and Möbius-Graph.
Figure 8 demonstrates the the query performance of Fargo on Commerce100M and Möbius-Graph on DBpedia1M, which is absent in Figure LABEL:fig:exp-qps due to either excessively high processing time or low recall. Specifically, on the Commerce100M dataset, Fargo achieves only 90% recall@100 and exhibits significantly slower performance. Möbius-Graph’s recall on DBpedia1M gets stuck at 78% because its vector space transformation approach, which, as discussed in §2, relies on assumptions that are frequently violated in real-world data, thereby impairing performance.
Figure 8. Query performance of ip-NSW, Fargo and Möbius-Graph.
Appendix E Empirical Verification of Theoretical Results.
Figure 9. A case study on the overlap of MIPS’s and NNS’s search paths with different configuration of $\mu$.
Figure 10. A case study on the overlap of overlap between k-MIPS solutions and the neighborhood around 1-MIPS solution.
Exp-1: Theory Verification (Q1). As detailed in §3, the key to equating $\mathsf{MIPS}$ with $\mathsf{NNS}$ on a proximity graph is finding an appropriate query-scaling factor $\mu$. We explore this property by conducting GNNS on an ideal $\mathsf{PSP}$ index. An ideal $\mathsf{PSP}$ can be obtained by executing Algorithm 2 with $L$ set to the cardinality of the base vectors and $R$ set to infinity. We perform this experiment on MNIST. Moreover, we evaluate Theorem 4 on synthetic datasets based on standard normal distribution, about the relationship between top-k $\mathsf{MIPS}$ solutions and the Euclidean neighborhood around top-1 $\mathsf{MIPS}$ solution.
Duality of NNS and MIPS on Proximity Graph. The case study is conducted by: (1) Randomly select a query and run GNNS with varying $\mu$ values: $\mu$ in $\{0.1,1,10,100,500\}$; (2) Trace node IDs visited during the search up to a maximum of $15$ steps, as MIPS usually reaches its solution within this limit. The results, as shown in Figure 9, demonstrate that the search paths of $\mathsf{MIPS}$ and $\mathsf{NNS}$ diverge initially but increasingly overlap as $\mu$ increases. At $\mu=100$, the search paths coincide completely. Furthermore, we assessed the average overlap ratio of search paths for NNS and MIPS, and recall@100 for NNS, across all queries on MNIST under different $\mu$ (as shown in Table 3). The result indicates that while precisely defining the lower bound of $\mu$ for every possible query is impractical, a sufficiently large $\mu$ can effectively unify the MIPS and NNS paths for all queries. Sub-optimal $\mu$ does not lead to collapsed results, yet yields reasonably good candidates due to highly overlapped paths.
Figure 11. The query process performance of various k of recall on MNIST, Text2image1M and DBpedia1M at 99% recall.
Overlap Between K-$\mathsf{MIPS}$ and Nearest Neighbors. Theorem 4 reveals a high confidence of the overlap between top-k $\mathsf{MIPS}$ solutions and the neighborhood around the top-1 $\mathsf{MIPS}$ solution. We evaluate how the overlap ratio changes with $s$ on standard Gaussian distribution datasets ($\mathcal{N}(0,1)$). The results shown in Figure 10 verify our analysis that after the top-k $\mathsf{MIPS}$ solutions are distributed around the top-1 $\mathsf{MIPS}$ solution’s Euclidean neighborhood with a high confidence. With higher dimensions and larger Euclidean distance range, the overlap gradually goes down, which aligns with intuition and the curse of dimensionality. In higher dimensional spaces, the relationship between vectors are more complicated. The duality of Euclidean and IP metric will degrade. Luckily, most real-world datasets, though with high dimension, lie on a very low-dimension manifold. For example, the intrinsic dimension of 960-dimensional GIST1M dataset is about 20+, making the search on it as simple as on a Gaussian random dataset of 20 dimensions.
Figure 12. Parameter sensitivity experimental results for key parameters and factors on Text2image1M and Text2image10M
Appendix F Parameter sensitivity experiments.
Figure 13. Experimental results of query performance using cosine similarity on the Music100 and Text2image1M datasets.
Parameter Sensitivity. Parameter tuning is often easier when patterns are predictable or when the parameter-performance relationship is smooth and convex. Our evaluations of PSP’s parameter sensitivity reveal several key consistent patterns of optimal configurations: (1) Parameters like $R$ (maximum out-degree) and $S$ (number of nodes added in EF) remain consistent across different dataset scales, as depicted in Figure 12. This consistency simplifies parameter tuning on a smaller subsets of data. (2) For the minimum angle between edges ($\alpha$), $60^{\circ}$ is optimal across all datasets, aligning with previous findings (Fu et al., 2021). (3) For other parameters like $Acc@K$ (accuracy of the $k$-NNG to initialize NSSG) and $L$ (candidate size for NSSG), larger values generally enhance accuracy but potentially increase processing time. Users can adjust these parameters based on specific performance requirements and computational constraints. In summary, parameter tuning for $\mathsf{PSP}$ indexing and searching is straightforward because of the predictable and consistent patterns observed across various datasets. The default configuration outlined in §5.1 ensures robust performance and is an alternative to parameter tuning, given $\mathsf{PSP}$’s demonstrated superiority across the evaluated datasets.
Appendix G Additional Scalability Experiments.
Scalability with Recall Quantity (k). To assess scalability relative to varying values of $k$, we adjusted $k$ from $\{1,10,20,50,100\}$ across the datasets MNIST, Text2Image1M, and DBpedia1M, each differing in cardinality and dimensionality. The results (Figure 11) suggest that both cardinality and dimensionality impact scalability in $k$. Higher cardinality and dimensionality both contribute to a faster increase in query processing time, aligning with expectations. Despite this, the scaling trend in query processing time remains practical due to the small basis of the time (lower than 1ms per query).
Scalability with Cosine Similarity Search. We perform experiments on the Music100 and Text2image1M datasets using cosine similarity search. We compared PSP with a state-of-the-art algorithm that first normalizes the dataset by scaling all data points to a norm of 1, constructs the graph using ip-NSW, and then applies cosine similarity search, referred to as normalized ip-NSW. In this case, cosine similarity, inner product, and nearest neighbor search are equivalent. As for PSP, We directly altered the metric to cosine similarity rather than normalizing the dataset. Figure 13 demonstrates that PSP yields favorable results when applying cosine similarity to unnormalized datasets. We also directly tested ip-NSW for cosine similarity search but encountered precision issues with Music100 (70%) and Text2image1M (89%). This stems from its tendency to connect nodes with large norms, which are less useful for cosine similarity. This highlights the effectiveness of PSP.
Appendix H Implications and Open Questions
Impact and Implication. This study represents the first attempt to align MIPS and NNS without altering the origin vector space, the first theoretical framework on search efficiency for graph-based $\mathsf{MIPS}$ methods, introducing a novel approach that could inspire a range of non-transformative methods. The theoretical insights provided have broad implications, extending beyond just graph-based vector retrieval to potentially influence other database and machine learning tasks.
Limitations. (1) The current study’s experiments may not fully capture performance variability across extremely skewed data distributions, particularly with outliers or distant clusters of varying norms. (2) Our theory does not specifically address redundant computation pruning tailored for the IP metric. Developing solutions for this issue could further enhance the efficiency of our method. (3) The research has not yet explored incremental indexing strategies.
Future Works. Building on the current study’s implications and its identified limitations, future research should focus on several key areas: (1) Further investigations should ascertain the adaptability and robustness of our methods across a broader array of heterogeneous datasets. (2) We consider developing acceleration techniques specifically tailored to the MIPS problem, backed by robust theoretical frameworks. (3) There is a need to explore incremental indexing techniques that can accommodate real-time updates to data without requiring a complete rebuilding of the index.